OpenStudy (anonymous):

solve for x: 27^(2x)=(1/9)^(x-3)

OpenStudy (anonymous):

trick is to write using the same base \[27=3^3\] and \[\frac{1}{9}=3^{-2}\]

OpenStudy (anonymous):

x= 3/4

OpenStudy (anonymous):

then you have \[(3^3)^{2x}=(3^{-3})^{x-3}\]

OpenStudy (anonymous):

x=3/4

OpenStudy (anonymous):

typo there solve \[(3^3)^{2x}=(3^{-2})^{x-3}\] i.e. \[6x=-2x+6\]

OpenStudy (anonymous):

i was 1st

OpenStudy (anonymous):

=3^6x=3^(-2x+6) 6x=-2x+6(after using exponent rule) 8x=6 x=3/4

OpenStudy (anonymous):

\[(3^3)^{2x}=\frac{1}{(3^2)^{(x-3)}}\] \[(3^2)^{3x}=\frac{1}{(3^2)^{(x-3)}}\] \[(3^2)^{3x}*(3^2)^{(x-3)}=1\] \[(3^2)^{3x+x-3}=(3^2)^0\] \[4x-3=0\] \[4x=3\] \[x=\frac{3}{4}\]

OpenStudy (anonymous):

Please let me know if any of the steps were unclear.